You are here

ENR CTIO Blanco 4-m 2015-09-01

Primary tabs

Date: 
Tuesday, September 1, 2015
Telescope: 
CTIO Blanco 4-m
Submitter: 
Kathy Vivas
Submitter Email: 
Observer Support: 
Hernan Tirado
Night Assistant: 
Alberto Alvarez
Also Present: 
David James (1st half), James Lasker (2nd half)
Problems: 

- DTS stopped during afternoon calibrations. Consequentely SISPI failed because it was
filling up all of it's available IBs. DTS was restarted in Tucson and things went back to normal.
Just in case, SISPI was restarted after calibrations

- At 23:08 local SISPI error --> FCM is now broken.
It looks like (from the GUI) that the compressed air supply has failed. On the Filter GUI of the ICS display,
both the Humidity and Supply fields are showing "{}; NOT OK" errors. Compressor was checked and it was opertional. The FCM role was restarted, and reconfigured, and the system came back up normally. It looks like the error was a lost connection with the FCM.

- At 23:45 local the Operator1 machine made and spontaneous shutdown, and it rebooted itself. Observations continued during this process.

- Around 3am local, all GUIs crashed at once and refreshing them did not bring them back (kept saying "Connecting to system1...."). We kill them and restarted them using the buttons in the top panel menu. The GUIs came back but not in their right places. We had to arrange them ourselves in the screens. At this point the Observer Console had "Level: user". We had to use the command releaseMasterConsole to gain master control. Observations continued and were monitored from observer2 during this process. No time lost.(thanks DJJ)

Comments: 

***DES (2nd half)***

*Observing Team*

_DES Observers:_ Kathy Vivas, James Lasker
_Telops:_ Hernan Tirado, Alberto Alvarez
_Remote Support:_

*Observing Plan*

# Community observer first half: David James (observer)
# Mid-night standards
# Turn on ObsTac
# Morning standards

*Conditions*

# Clear, beautiful night. Seeing < 1 arcsec during most of the night. But.... very bright moon

*Accomplishments*

# 2 standard star fields: SDSSJ2300-0000, C2620/HST
# Survey exposures mostly in Y (a few ones in z at the start of the shift)
# GPS check (OK), aTmCam check (OK)
# 2 standard star fields: SDSSJ0100-0000, SDSSJ0320-0000

Conditions: 
General Conditions
Comment: 
None.
1st Quarter
Clouds: 
Clear
Wind: 
Low
Seeing Value: 
1.00
Seeing Variability: 
0.2
Seeing Source: 
The Telescope I used
Sky : 
0
2nd Quarter
Clouds: 
Clear
Wind: 
Low
Seeing Value: 
1.00
Seeing Variability: 
0.2
Seeing Source: 
The Telescope I used
Sky: 
0
3rd Quarter
Clouds: 
Clear
Wind: 
Low
Seeing Value: 
0.80
Seeing Variability: 
0.2
Seeing Source: 
The Telescope I used
Sky: 
0
4th Quarter
Clouds: 
Clear
Seeing Value: 
1.20
Seeing Variability: 
0.2
Seeing Source: 
CTIO Seeing monitor
Sky: 
0
Program: 
Proposal: 
2015A-0314
Target of Oportunity: 
No
Time Observed: 
5.0
Time Engineering: 
0.0
Time lost technical: 
0.0
Time lost weather: 
0.0
Time lost other: 
0.0
Description: 
Catching Supernovae in the Act with KEGS (Kepler Extra-Galactic Survey) and DECam
Primary Investigator: 
B. Tucker
Institution: 
U.Berkeley
Instrument: 
CTIO 4m DECam
Organization: 
NOAO
Total Program Time: 
5
Proposal: 
2012B-0001
Target of Oportunity: 
No
Time Observed: 
5.0
Time Engineering: 
0.0
Time lost technical: 
0.0
Time lost weather: 
0.0
Time lost other: 
0.0
Description: 
DES
Primary Investigator: 
Frieman
Institution: 
Fermilab
Instrument: 
CTIO 4m DECam
Organization: 
NOAO
Total Program Time: 
5
Total Times Time Observed Time Engineering Time lost technical Time lost weather Time lost other Total Program Time
2015-09-01 10.0 0.0 0.0 0.0 0.0 10